OpenStudy (anonymous):

Solve for x. 5a/x = 5b/x-1

OpenStudy (ankit042):

cross multiply 5a(x-1) = 5b(x NOw simplify

OpenStudy (compassionate):

\[\frac{ 5a }{ x } = \frac{ 5b }{ x-1 }\] The first step is cross-multiplying. This eliminates the fraction and you're left with. 5ax - 5a = 5bx